# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ 8 Conclusion
In this article, we evaluate the performance of Gaussian variational inference based on the weighted Fisher divergence by focusing on the FD and SD. First, we consider the mean-field assumption for both Gaussian and non-Gaussian targets. We demonstrate that FD and SD tend to underestimate the posterior variance more severely than KLD for Gaussian targets in this context. SD has the potential to capture the posterior mode more accurately than FD and KLD for skewed targets, but underestimates the variance more severely.
@@
Next, we consider high-dimensional hierarchical models for which posterior conditional independence structure can be captured using a sparse precision matrix in the Gaussian variational approximation. To impose sparsity on the Cholesky factor of the precision matrix, we consider optimization based on SGD and propose two approaches based on the reparametrization trick and a batch approximation of the objective.
@@
The reparametrization trick yields unbiased gradient estimates but involves a Hessian matrix, which can be computationally expensive and increases variability in the gradients, leading to reduced stability and slow convergence. To address these issues, we introduce an alternative that minimizes a biased estimate of the FD and SD using a random batch of samples at each iteration. This eliminates reliance on the Hessian matrix, improving stability. This approach can also be interpreted as optimizing a new objective that iteratively improves the match between gradients of the posterior and variational density, at sample points that shift gradually towards high probability regions of the posterior. We evaluate the behavior of this new objective under the mean-field assumption for Gaussian targets and show that it alleviates the variational collapse issue faced previously by SD.
@@
The proposed methods are compared to KLD and BaM in applications involving logistic regression, GLMMs and stochastic volatility models. Extensive experiments reveal that FDr and SDr converge very slowly, often to suboptimal variational approximations. FDb and SDb provide substantial improvements over FDr and SDr, with SDb having superior performance in terms of convergence rate and quality of variational approximation attained. BaM, which relies on closed-form updates and hence requires fewer iterations to converge, is very effective for logistic regression. However, it is much less efficient than KLD for GLMMs and SSMs, and its performance gradually worsens as the dimension increases, eventually failing to converge. SDb has an advantage over BaM in high dimensions as it is able to impose sparsity on the precision matrix, remains feasible computationally and is more stable and less sensitive to poor initialization. SDb can also capture posterior modes more accurately than KLD although it is more prone to variance underestimation.
@@
Several avenues for future research remain open. Our analysis focused on Gaussian variational approximations for FD and SD, and it will be valuable to investigate the performance of FD and SD under more flexible variational families. While we have used SGD for optimization, the choice of optimizer and associated hyperparameters significantly influences convergence behavior, and it will be useful to explore alternative optimization techniques that do not rely on SGD. Our findings also highlight the potential of the batch approximated SD, and it will be interesting to further investigate its properties in other contexts.
@@
S1 Univariate non-Gaussian target
We begin by deriving some key expressions that will be used throughout our analysis of non-Gaussian target distributions. We assume the variational approximation $q(\theta)=\phi(\theta|\mu,\sigma^{2})$. Its log-density and gradient are
@@
$\nabla_{\theta}\log q(\theta)$ $=-\frac{\theta-\mu}{\sigma^{2}}.$
The expectation of $\log q(\theta)$ with respect to $q(\theta)$ is
@@
S1.1 Student’s $t$
The log-density and its gradient for $\theta\sim t(\nu)$ are
$\log p(\theta,y)$ $=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log(\pi\nu)$
@@
$\nabla\log p(\theta,y)$ $=-(\nu+1)\frac{\theta}{\nu+\theta^{2}},$
where $\nu$ is the degree of freedom. The evidence lower bound for KLD is
${\mathcal{L}}$ $={\text{E}}_{q}\left\{\log p(\theta,y)-\log q(\theta)\right\}$
$=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log(\pi\nu)-\log\left\{\Gamma\left(\frac{\nu}{2}\right)\right\}$
$\quad-\frac{\nu+1}{2}{\text{E}}_{q}\log\left(1+\frac{\theta^{2}}{\nu}\right)+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
@@
$F(q\|p)$ $={\text{E}}_{q}\{\|\nabla_{\theta}\log p(\theta,y)-\nabla_{\theta}\log q(\theta)\|^{2}\}$
@@
S1.2 Log transformed inverse Gamma
For the log transformed inverse gamma, let $a_{1}=a_{0}+n/2$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. Then its log-density and gradient are given by
@@
$=-\frac{n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})-a_{1}\mu$
$\quad-\exp\left(\frac{\sigma^{2}}{2}-\mu\right)b_{1}+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
For the Fisher divergence, we evaluate the expectation of the squared norm of the gradient difference,
@@
$\quad+2b_{1}{\text{E}}_{q}\left\{\exp(-\theta)\frac{\theta-\mu}{\sigma^{2}}\right\}.$
Note that ${\text{E}}_{q}\{\exp(-a\theta)\}=\exp(a^{2}\sigma^{2}/2-a\mu)$ for any constant $a\in\mathbb{R}$. Thus,
$F(q\|p)$ $=a_{1}^{2}+b_{1}^{2}\exp(2\sigma^{2}-2\mu)$
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)+1/\sigma^{2}.$
We have
@@
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)\}+1.$
$\nabla_{\mu}S(q\|p)$ $=2b_{1}\sigma^{2}\exp(\sigma^{2}/2-\mu)\{(a_{1}+1)$
$\quad-b_{1}\exp(3\sigma^{2}/2-\mu)\}=0$
$\therefore\mu$ $=\log\frac{b_{1}}{a_{1}+1}+\frac{3\sigma^{2}}{2}.$
@@
$S(q\|p)$ $=a_{1}^{2}\sigma^{2}-(a_{1}+1)^{2}\sigma^{2}\exp(-\sigma^{2})+1.$
$\nabla_{\sigma^{2}}S(q\|p)$ $=a_{1}^{2}+(a_{1}+1)^{2}\exp(-\sigma^{2})(\sigma^{2}-1).$
S1.3 Skew normal
@@
$p(\theta,y)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}.$
@@
$\log p(\theta,y)$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{(\theta-m)^{2}}{2t^{2}}$
@@
$\nabla_{\theta}\log p(\theta,y)$ $=-\frac{\theta-m}{t^{2}}+\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}.$
Taking the expectation of $\log p(\theta,y)$ with respect to $q(\theta)$,
${\text{E}}_{q}\left\{\log p(\theta,y)\right\}$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
@@
${\mathcal{L}}$ $={\text{E}}_{q}\left\{\log p(\theta,y)-\log q(\theta)\right\}$
$=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
$\quad+{\text{E}}_{q}\log[\Phi\{\lambda(\theta-m)\}]+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is given by
$F(q\|p)={\text{E}}_{q}\{\|\nabla_{\theta}\log p(\theta,y)-\nabla_{\theta}\log q(\theta)\|^{2}\}$
@@
S2 SGD based on reparametrization trick
@@
$d\theta=d\mu\end{aligned}\quad\text{and}\quad\begin{aligned} d\theta=-T^{-\top}(dT^{\top})T^{-\top}z.\end{aligned}$

### v2

@@ 8 Conclusion
In this article, we evaluate the performance of Gaussian variational inference based on the weighted Fisher divergence by focusing on the FD and SD. First, we consider the mean-field assumption for Gaussian and non-Gaussian targets. We demonstrate that FD and SD tend to underestimate the posterior variance more severely than KLD, and SD can capture the posterior mode more accurately than FD and KLD for skewed targets.
@@
Next, we consider high-dimensional hierarchical models whose posterior conditional independence structure can be captured using a sparse precision matrix in the Gaussian variational approximation. To impose sparsity on the Cholesky factor of the precision matrix, we consider optimization based on SGD and propose two approaches based on the reparametrization trick and a batch approximation of the objective.
@@
The reparametrization trick yields unbiased gradient estimates but involves a Hessian matrix, which is computationally expensive and increases variability in the gradients, leading to reduced stability and slow convergence. To address these issues, we introduce an alternative that minimizes a biased estimate of the FD and SD computed using a random batch of samples at each iteration. This eliminates reliance on the Hessian and improves stability. This approach can also be interpreted as optimizing a new objective, that iteratively improves the match between gradients of the posterior and variational density, at sample points that shift gradually towards regions of high posterior probability. While the general convergence of FDb and SDb remains as an open problem, we make some contribution in this direction by proving the convergence of SDb in the special case of a Gaussian target with infinite batch size, using natural gradients updates with a constant stepsize. We also evaluate the behavior of this new objective under the mean-field assumption for Gaussian targets and show that it alleviates the variational collapse issue faced previously by SD.
@@
The proposed methods are compared to KLD, (p)BaM in applications involving logistic regression, GLMMs and stochastic volatility models. Extensive experiments reveal that FDr and SDr converge very slowly, often to suboptimal variational approximations. FDb and SDb provide substantial improvements over FDr and SDr, with SDb having superior performance in terms of convergence rate and accuracy. (p)BaM, which relies on closed-form updates and hence requires fewer iterations to converge, is very effective for logistic regression. However, it is less efficient than KLD for GLMMs and stochastic volatility models, and its performance gradually worsens as the dimension increases, eventually failing to converge. SDb has an advantage over (p)BaM in high dimensions as it can impose sparsity on the precision matrix, remains feasible computationally and is more stable and less sensitive to poor initialization. SDb can also capture posterior modes more accurately than KLD but is more prone to variance underestimation.
@@
There are several avenues for future research. While this work has focused primarily on two variants of the weighted Fisher divergence (FD and SD), it will be valuable to investigate other variants. Besides Gaussian variational approximations, it is also of interest to investigate the performance of FD and SD under more flexible variational families. While we have used SGD for optimization, the choice of optimizer and associated hyperparameters significantly influences convergence behavior, and it is useful to explore alternative optimization techniques based on natural gradients or which do not rely on SGD. Our findings highlight the potential of the batch approximated SD, and its properties can be investigated further in other contexts. Finally, proving the convergence of our batch-approximated methods in the practical setting with non-Gaussian targets and finite batches remains as an open problem.
9 Acknowledgment
We would like to thank the Editor, Associate Editor and three referees for their comments and helpful suggestions which have improved this manuscript greatly.
@@ 9 Acknowledgment
S1 Proof of Lemma 1
The $M$-weighted Fisher divergence is
$=\text{\rm tr}\{(\Sigma^{-1}-\Lambda)M(\Sigma^{-1}-\Lambda)\Sigma\}$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu)$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu).$
The final result arises from $\text{\rm tr}(AB)=\sum_{i=1}^{d}A_{ii}B_{ii}$ if $A$ is a diagonal matrix.
S2 Proof of Theorem 2
First we present Lemma S1, which is required in the proof of Theorem 2.
Lemma S1.
Let $\theta\sim{\text{N}}(\mu,\Sigma)$. If $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ is integrable and is an odd function of $(\theta-\mu)$ in that $f(\mu-\theta)=-f(\theta-\mu)$, then ${\text{E}}_{\theta\sim{\text{N}}(\mu,\Sigma)}[f(\theta-\mu)]=0$.
Proof.
Let $\theta^{\prime}=\theta-\mu$ so that $\theta^{\prime}\sim{\text{N}}(0,\Sigma)$. Then $f(\theta^{\prime})=f(\theta-\mu)=-f(\mu-\theta)=-f(-\theta^{\prime})$.
${\text{E}}_{\theta\sim{\text{N}}(\mu,\Sigma)}$ $[f(\theta-\mu)]={\text{E}}_{\theta^{\prime}\sim{\text{N}}(0,\Sigma)}[f(\theta^{\prime})]$
$=\int_{\mathbb{R}^{d}}f(\theta^{\prime})\phi(\theta^{\prime}\mid 0,\Sigma)d\theta^{\prime}$
$=\int_{\mathbb{R}^{d}}f(-\theta^{\prime})\phi(-\theta^{\prime}\mid 0,\Sigma)d\theta^{\prime}$
$=\int_{\mathbb{R}^{d}}\{-f(\theta^{\prime})\}\phi(\theta^{\prime}\mid 0,\Sigma)d\theta^{\prime}$
$=-{\text{E}}_{\theta^{\prime}\sim{\text{N}}(0,\Sigma)}[f(\theta^{\prime})]$
$=-{\text{E}}_{\theta\sim{\text{N}}(\mu,\Sigma)}[f(\theta-\mu)]$
Thus, ${\text{E}}_{\theta\sim{\text{N}}(\mu,\Sigma)}[f(\theta-\mu)]=-{\text{E}}_{\theta\sim{\text{N}}(\mu,\Sigma)}[f(\theta-\mu)]=0$. ∎
For $q(\theta)={\text{N}}(\mu,\Sigma)$, $\nabla_{\theta}\log q(\theta)=-\Sigma^{-1}(\theta-\mu)$. For the multivariate Student’s $t$ distribution, let
$\delta(\theta)=(\theta-m)^{\top}S^{-1}(\theta-m),\quad w(\theta)=\frac{\nu+d}{\nu+\delta(\theta)}.$
Then
$\nabla_{\theta}\log p(y,\theta)$ $=-w(\theta)S^{-1}(\theta-m),$
$H_{p}(\theta)$ $=\nabla_{\theta}^{2}\log p(y,\theta)=-w(\theta)S^{-1}$
$\quad+\frac{2w(\theta)}{\nu+\delta(\theta)}S^{-1}(\theta-m)(\theta-m)^{\top}S^{-1}.$
For the KLD, the evidence lower bound ${\mathcal{L}}={\text{E}}_{q}[\log p(y,\theta)-\log q(\theta)]$. Since ${\text{E}}_{q}[\log q(\theta)]$ is independent of $\mu$, we only need to focus on the first term. By applying the reparametrization trick described in Section 5 of the manuscript,
$\nabla_{\mu}\mathcal{L}$ $={\text{E}}_{q}[\nabla_{\theta}\log p(y,\theta)]=-{\text{E}}_{q}[w(\theta)S^{-1}(\theta-m)].$
Here $w(\theta)$ is even in $(\theta-m)$, while $(\theta-m)$ is odd, so the integrand is odd in $(\theta-m)$. Lemma S1 implies that $\nabla_{\mu}\mathcal{L}=-{\text{E}}_{q}[w(\theta)S^{-1}(\theta-m)]=0$ at $\mu=m$.
For the FD, from Section 5 of the manuscript,
$\nabla_{\mu}F(q\|p)=2{\text{E}}_{q}[H_{p}(\theta)(\nabla_{\theta}\log p(y,\theta)-\nabla_{\theta}\log q(\theta))].$
Note that $H_{p}(\theta)$ is even in $(\theta-m)$. Moreover,
$\nabla_{\theta}\log p(y,\theta)-\nabla_{\theta}\log q(\theta)\\ =-w(\theta)S^{-1}(\theta-m)+\Sigma^{-1}(\theta-m),$
which is odd in $(\theta-m)$. Thus the integrand in $\nabla_{\mu}F(q\|p)$ is odd, and by Lemma S1, $\nabla_{\mu}F(q\|p)=0$ at $\mu=m$.
For the SD, from Section 5 of the manuscript,
$\nabla_{\mu}S(q\|p)=2{\text{E}}_{q}[H_{p}(\theta)\Sigma(\nabla_{\theta}\log p(y,\theta)-\nabla_{\theta}\log q(\theta))].$
The same argument as for FD shows that the integrand is odd in $(\theta-m)$, and hence $\nabla_{\mu}S(q\|p)=0$ at $\mu=m$.
Hence $\mu=m$ is a stationary point for all three divergences.
S3 Univariate non-Gaussian target
We begin by deriving some key expressions that are used throughout our analysis of non-Gaussian target distributions. Suppose the variational approximation $q(\theta)=\phi(\theta|\mu,\sigma^{2})$. We have
@@
$\nabla_{\theta}\log q(\theta)$ $=-\frac{\theta-\mu}{\sigma^{2}},$
@@
For univariate densities, $S(q\|p)=\sigma^{2}F(q\|p)$, so only the expression of FD is presented.
S3.1 Student’s $t$
When the target is the univariate Student’s $t$, such that $\theta\sim t(\nu)$, where $\nu$ is the degree of freedom,
$\log p(y,\theta)$ $=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log(\pi\nu)$
@@
$\nabla\log p(y,\theta)$ $=-\frac{(\nu+1)\theta}{\nu+\theta^{2}}.$
The evidence lower bound for KLD is
$=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log\left\{\Gamma^{2}\left(\frac{\nu}{2}\right)\nu\right\}$
@@
S3.2 Log transformed inverse Gamma
For the log transformed inverse gamma target, recall that $a_{1}=a_{0}+n/2>1/2$ since $n\geq 1$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. Then
@@
Setting $\nabla_{\theta}\log p(y,\theta)=0$, the true posterior mode $m_{*}=\log(b_{1}/a_{1})$. Since $\exp(-\theta)\mid y\sim\text{Gamma}(a_{1},b_{1})$, the true posterior mean and variance are given by $\mu_{*}={\text{E}}(\theta\mid y)=\log(b_{1})-\psi(a_{1})$ and $\sigma_{*}^{2}=\text{\rm Var}(\theta\mid y)=\psi_{1}(a_{1})$ respectively (pg. 33, Hall and Oakes, 2024). As $\psi(a_{1})<\log(a_{1})$ (Alzer, 1997), $\mu_{*}>m_{*}$ and the true posterior is right skewed.
First, we find the optimal variational parameters $(\hat{\mu}_{\text{KL}},\hat{\sigma}^{2}_{\text{KL}})$ that maximize the evidence lower bound for the KLD. We have
@@
$=\frac{1-n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})-a_{1}\mu$
$\quad-b_{1}\exp\left(\frac{\sigma^{2}}{2}-\mu\right)+\frac{1}{2}\log(\sigma^{2})+\frac{1}{2}.$
Setting
$\nabla_{\mu}{\mathcal{L}}$ $=-a_{1}+b_{1}\exp(\sigma^{2}/2-\mu)=0,$
$\nabla_{\sigma^{2}}{\mathcal{L}}$ $=-\frac{b_{1}}{2}\exp(\sigma^{2}/2-\mu)+\frac{1}{2\sigma^{2}}=0.$
and solving simultaneously gives the global maximum at
$\hat{\mu}_{\text{\rm KL}}$ $=\log\!\Big(\frac{b_{1}}{a_{1}}\Big)+\frac{1}{2a_{1}},\quad\hat{\sigma}^{2}_{\text{\rm KL}}=\frac{1}{a_{1}}.$
Now, we find the optimal variational parameters $(\hat{\mu}_{\text{F}},\hat{\sigma}^{2}_{\text{F}})$ and $(\hat{\mu}_{\text{S}},\hat{\sigma}^{2}_{\text{S}})$ that minimize the FD and SD respectively. For the FD,
@@
$\quad+2b_{1}{\text{E}}_{q}\left\{\exp(-\theta)\frac{\theta-\mu}{\sigma^{2}}\right\}$
$=a_{1}^{2}+b_{1}^{2}\exp(2\sigma^{2}-2\mu)$
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)+1/\sigma^{2},$
since ${\text{E}}_{q}\{\exp(-a\theta)\}=\exp(a^{2}\sigma^{2}/2-a\mu)$ for any constant $a\in\mathbb{R}$. It follows that
@@
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)\}+1,$
$\nabla_{\mu}S(q\|p)$ $=2b_{1}\sigma^{2}\exp(\sigma^{2}/2-\mu)\{a_{1}+1$
$\quad-b_{1}\exp(3\sigma^{2}/2-\mu)\}.$
Note that $\nabla_{\mu}S(q\|p)=\sigma^{2}\nabla_{\mu}F(q\|p)$. Therefore, setting $\nabla_{\mu}S(q\|p)=0$ and $\nabla_{\mu}F(q\|p)=0$ both lead to the same condition,
$\mu=\log\frac{b_{1}}{a_{1}+1}+\frac{3\sigma^{2}}{2}.$
@@
$F(\sigma^{2})$ $=F(q\|p)|_{\mu=\hat{\mu}_{F}}$
$=a_{1}^{2}-(a_{1}+1)^{2}\exp(-\sigma^{2})+\frac{1}{\sigma^{2}}.$
$S(\sigma^{2})$ $=S(q\|p)|_{\mu=\hat{\mu}_{S}}$
$=a_{1}^{2}\sigma^{2}-(a_{1}+1)^{2}\sigma^{2}\exp(-\sigma^{2})+1.$
Setting
$F^{\prime}(\sigma^{2})$ $=(a_{1}+1)^{2}\exp(-\sigma^{2})-\frac{1}{\sigma^{4}}=0,$
$S^{\prime}(\sigma^{2})$ $=a_{1}^{2}+(a_{1}+1)^{2}(\sigma^{2}-1)\exp(-\sigma^{2})=0,$
we obtain
$\hat{\sigma}^{2}_{\text{F}}$ $=-2W_{0}\left(-\frac{1}{2(a_{1}+1)}\right),$
$\hat{\sigma}^{2}_{\text{S}}$ $=1-W_{0}\left(\frac{{\rm e}a_{1}^{2}}{(a_{1}+1)^{2}}\right),$
where $W_{0}$ is the principal branch of the Lambert W function (Corless et al., 1996). The Lambert W function yields the solution to the equation $z\exp(z)=a$, such that $z=W_{0}(a)\geq 0$ if $a\geq 0$, and either $z=W_{0}(a)\in[-1,0)$ or $z=W_{-1}(a)\leq-1$ if ${\rm e}^{-1}\leq a<0$. For SD, the argument $0<{{\rm e}a_{1}^{2}}/{(a_{1}+1)^{2}}<{\rm e}$ and hence $S(\sigma^{2})$ has a global minimum at $\hat{\sigma}^{2}_{\text{S}}\in(0,1)$. For FD, it can be verified that ${\rm e}^{-1}<-1/\{2(a_{1}+1)\}<0$ and hence $F(\sigma^{2})$ has two stationary points, one in $(0,2)$ and the other in $(2,\infty)$. As $\lim_{\sigma^{2}\rightarrow 0^{+}}F(\sigma^{2})=+\infty$ and $\lim_{\sigma^{2}\rightarrow+\infty}F(\sigma^{2})=a_{1}^{2}$, the global minimum occurs in $(0,2)$ and is given by the principal branch $W_{0}(\cdot)$. Plots of $F(\sigma^{2})$ and $S(\sigma^{2})$ are given in Fig S1. It follows that
$\hat{\mu}_{\text{F}}$ $=\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{F}}}{2},$ (S1)
$\hat{\mu}_{\text{S}}$ $=\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{S}}}{2}.$
Figure S1: Plots of $F(\sigma^{2})$ and $S(\sigma^{2})$ when $a_{1}=0.501$.
Before proving Theorem 3, we require some intermediate results, which are summarized in Lemma S2.
Lemma S2.
  1. i.
    Let $h(x)=x/2-\log(x)-\log(a_{1}+1)$. Then $h(1/a_{1})<0$ for $a_{1}>1/2$ .
  2. ii.
    $\psi_{1}(x)>1/x$ for $x>0$.
  3. iii.
    $S^{\prime}(c)<0$ where $c=(2/3)\log(1+1/a_{1})$.
  4. iv.
    $(2/3)\log(1+1/a_{1})+1/(3a_{1})>b$ where $b=(2/3)/(a_{1}+1/2)+1/(3a_{1})$.
  5. v.
    $f(a_{1})=\log(a_{1}+1)+\log(b)-b/2>0$ for $a_{1}>1/2$ where $b$ is defined in iv.
Proof.
For i, we want to show that $h(1/a_{1})=\log(a_{1})+1/(2a_{1})-\log(a_{1}+1)<0$. Let $g(x)=\log x+1/(2x)-\log(x+1)$. Then $g^{\prime}(x)=(x-1)/\{2x^{2}(x+1)\}$. Hence $g$ is decreasing on $(1/2,1]$ and increasing on $1,\infty)$, with g(1/2)=2/3g(1/2)=-2/3 and $\lim_{x\to+\infty}g(x)=0^{-}$. Thus $g(x)<0$ for all $x>\frac{1}{2}$, and hence $h(1/a_{1})<0$.
For ii, we can write $\psi_{1}(x)=\sum_{n=0}^{\infty}1/(n+x)^{2}$ and $1/x=\sum_{n=0}^{\infty}\{1/(n+x)-1/(n+x+1)\}$. For $x>0$,
$\psi_{1}(x)-\frac{1}{x}=\sum_{n=0}^{\infty}\frac{1}{n+x}\left(\frac{1}{n+x}-\frac{1}{n+x+1}\right)>0.$
For iii, $S^{\prime}(c)=a_{1}^{2}+(a_{1}+1)^{2}(c-1)\exp(-c)<0$ is equivalent to $\log(1-c)+2c>0$ since $\log\{a_{1}/(a_{1}+1)\}=-3c/2$. Let $k(x)=\log(1-x)+2x$ for $0<x<1$. Then $k^{\prime}(x)=(1-2x)/(1-x)$. Thus $k(x)$ increases from $(0,0)$, reaches a maximum at $(1/2,\log(1/2)+1)$ and decreases to -$\infty$ with an asymptote at $x=1$. Since $0<c<(2/3)\log(3)$, and $k(2/3\log 3)=0.147>0$, we conclude that $k(c)>0$ and hence $S^{\prime}(c)<0$.
For iv, we use the inequality $\log(1+x)>x/(1+x/2)$ for $x>0$ from Love (1980), which implies that $\log(1+1/a_{1})>1/(a_{1}+1/2)$.
For v, as $b=(a_{1}+1/6)/\{a_{1}(a_{1}+1/2)\}$
$f^{\prime}(a_{1})=\frac{1}{a_{1}+1}+\frac{1}{a_{1}+1/6}-\frac{1}{a_{1}}-\frac{1}{a_{1}+1/2}+\frac{1}{6a_{1}^{2}}\\ -\frac{1}{3(a_{1}+1/2)^{2}}=-\frac{120a_{1}^{4}+100a_{1}^{3}+6a_{1}^{2}-5a_{1}-1}{6a_{1}^{2}(a_{1}+1)(2a_{1}+1)^{2}(6a_{1}+1)}.$
It can be verified that $f^{\prime}(a_{1})<0$ and hence $f(a_{1})$ is strictly decreasing for $a_{1}>1/2$. Moreover $f(0.5)=\log(2)-2/3>0$ and $\lim_{a_{1}\to+\infty}f(a_{1})=0^{+}$. Hence $f(a_{1})>0$ for $a_{1}>1/2$. ∎
Proof of Theorem 3.
First, we establish the ordering for the variance parameters. At the global minimum of FD, we have $F^{\prime}(\hat{\sigma}^{2}_{\text{F}})=0$. As $S(\sigma^{2})=\sigma^{2}F(\sigma^{2})$,
$S^{\prime}(\hat{\sigma}^{2}_{\text{F}})=F(\hat{\sigma}^{2}_{\text{F}})+\hat{\sigma}^{2}_{\text{F}}F^{\prime}(\hat{\sigma}^{2}_{\text{F}})=F(\hat{\sigma}^{2}_{\text{F}})>0.$
The strict inequality above holds because $F(q\|p)=0$ if and only if $q(\theta)=p(\theta\mid y)$ almost everywhere. However, this is not true as $q(\theta)$ is symmetric while $p(\theta\mid y)$ is right skewed. Since $S(\sigma^{2})$ has a global minimum and $S^{\prime}(\hat{\sigma}^{2}_{\text{F}})>0$, its minimum must occur strictly before $\hat{\sigma}^{2}_{\text{F}}$, and hence
$\hat{\sigma}^{2}_{\text{S}}<\hat{\sigma}^{2}_{\text{F}}.$
Next, we want to show that $0<\hat{\sigma}^{2}_{\text{F}}<\hat{\sigma}^{2}_{\text{\rm KL}}=1/a_{1}$ ¡ 2. By setting $F^{\prime}(\sigma^{2})=0$, observe that $h(\sigma^{2})=\sigma^{2}/2-\log(\sigma^{2})-\log(a_{1}+1)$ is strictly decreasing and has a single root $\hat{\sigma}^{2}_{\text{F}}$ in $(0,2)$. Moreover, $h(1/a_{1})<0$ from Lemma S2i. Hence $\hat{\sigma}^{2}_{\text{F}}<\hat{\sigma}^{2}_{\text{\rm KL}}$. Finally, $\hat{\sigma}^{2}_{\text{\rm KL}}=1/a_{1}<\psi_{1}(a_{1})=\sigma_{*}^{2}$ from Lemma S2ii. Hence,
$\hat{\sigma}^{2}_{\text{S}}<\hat{\sigma}^{2}_{\text{F}}<\hat{\sigma}^{2}_{\text{\rm KL}}<\sigma_{*}^{2}.$
Next, we establish the ordering for the mean parameters. First, $\hat{\sigma}^{2}_{\text{S}}<\hat{\sigma}^{2}_{\text{F}}$ implies that $\hat{\mu}_{\text{S}}<\hat{\mu}_{\text{F}}$ from (S1). Next,
$\mu_{*}-\hat{\mu}_{\text{\rm KL}}$ $=\{\log b_{1}-\psi(a_{1})\}-\left(\log\frac{b_{1}}{a_{1}}+\frac{1}{2a_{1}}\right)$
$=\log a_{1}-\psi(a_{1})-\frac{1}{2a_{1}}>0$
for $a_{1}>0$, based on a result from (pg 374, Alzer, 1997). Hence $\hat{\mu}_{\text{\rm KL}}<\mu_{*}$. We want to show that
$\hat{\mu}_{\text{S}}-m_{*}$ $=\left(\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{S}}}{2}\right)-\log\frac{b_{1}}{a_{1}}$
$=\frac{3}{2}\hat{\sigma}^{2}_{\text{S}}-\log\left(1+\frac{1}{a_{1}}\right)>0.$
This is equivalent to showing that $\hat{\sigma}^{2}_{\text{S}}>c$, where $0<c=(2/3)\log(1+1/a_{1})<(2/3)\log(3)\approx 0.732$. Recall that $S(\sigma^{2})$ has a global minimum at $\hat{\sigma}^{2}_{\text{S}}$. Thus, it suffices to show that $S^{\prime}(c)<0$, which holds from Lemma S2iii.
It remains to show that $\hat{\mu}_{\text{F}}<\hat{\mu}_{\text{\rm KL}}$. We have
$\hat{\mu}_{\text{\rm KL}}-\hat{\mu}_{\text{F}}$ $=\left(\frac{1}{2a_{1}}+\log\frac{b_{1}}{a_{1}}\right)-\left(\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{F}}}{2}\right)$
$=\log\left(\frac{a_{1}+1}{a_{1}}\right)+\frac{1}{2a_{1}}-\frac{3}{2}\hat{\sigma}^{2}_{\text{F}}.$
Hence, our goal is to show that $\hat{\sigma}^{2}_{\text{F}}<(2/3)\log(1+1/a_{1})+1/(3a_{1})$. We split the proof into two parts by showing that $\hat{\sigma}^{2}_{\text{F}}<b$ and $b\leq(2/3)\log(1+1/a_{1})+1/(3a_{1})$, where $0<b=(2/3)/(a_{1}+1/2)+1/(3a_{1})<4/3$. The second part of the proof is given in Lemma S2iv. For the first part of the proof, recall that $F(\sigma^{2})$ has a single global minimum in $(0,2)$. Hence it suffices to show that $F^{\prime}(b)>0$ or equivalently that $\log(a_{1}+1)+\log(b)-b/2>0$. This is true from Lemma S2v.
Therefore, we have $m_{*}<\hat{\mu}_{\text{S}}<\hat{\mu}_{\text{F}}<\hat{\mu}_{\text{\rm KL}}<\mu_{*}$. ∎
S3.3 Skew normal
@@
$p(y,\theta)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}.$
@@
$\log p(y,\theta)$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{(\theta-m)^{2}}{2t^{2}}$
@@
$\nabla_{\theta}\log p(y,\theta)$ $=-\frac{\theta-m}{t^{2}}+\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}.$
${\text{E}}_{q}\left\{\log p(y,\theta)\right\}$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
@@
$=\log 2-\log(t)-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
The FD is given by
@@
S4 SGD based on reparametrization trick
@@
$\begin{aligned} d\theta=d\mu\end{aligned}\quad\text{and}\quad\begin{aligned} d\theta=-T^{-\top}(dT^{\top})T^{-\top}z.\end{aligned}$